# The asymptotic dimension of the grand arc graph is infinite

### v2

@@ 4.1. Admissible combinatorial models
A cleanly marked arc and curve system $\omega$ on $\Omega$ is the union of an arc system $\alpha$ and a marking $\mu$ on $\Omega$ such that:
@@
    $\alpha,\operatorname{base}\mu$ are disjoint, and
@@
    the maximal submarking $\mu^{\prime}$ with only non-empty transversals is complete and clean in each component of $\Omega\setminus(\omega\setminus\mu^{\prime})$ that it intersects.
@@
A marking which satisfies the above is locally clean.
Let $\operatorname{base}\omega=\alpha\cup\operatorname{base}\mu$ denote the underlying arc and curve system. For a component $a\in\operatorname{base}\omega$, let $\operatorname{trans}_{\omega}(a)\mathrel{\mathop{:}}=\operatorname{trans}_{\mu}(a)$ if $a\in\operatorname{base}\mu$, else $\varnothing$ if $a\in\alpha$. We define the geometric intersection of $\omega=\alpha\cup\mu$ and $\omega^{\prime}=\alpha^{\prime}\cup\mu^{\prime}$ to be
$i(\omega,\omega^{\prime})\mathrel{\mathop{:}}=i(\operatorname{base}\omega,\operatorname{base}\omega^{\prime})+\sum_{a\in\operatorname{base}\mu\cup\operatorname{base}\mu^{\prime}}\operatorname{diam}_{\mathcal{C}(a)}(\pi_{a}\omega\cup\pi_{a}\omega^{\prime})$
where $\pi_{a}(\omega)\mathrel{\mathop{:}}=\pi_{a}\alpha\cup\pi_{a}\mu$ and likewise for $\omega^{\prime}$. An essential, non-pants subsurface intersects $\omega$ if and only if it intersects $\alpha$ or $\mu$.
Notation.
Let $\mathcal{MS}(\Omega)$ denote the set of cleanly marked arc and curve systems on $\Omega$.
@@
A multiarc and curve graph $\mathcal{A}$ on $\Omega$ is a simplicial graph whose vertices are arc and curve systems on $\Omega$. Likewise, a marking graph $\mathcal{M}$ on $\Omega$ is a simplicial graph whose vertices are locally clean markings. Most generally, a combinatorial model for $\Omega$ is a simplicial graph whose vertices are cleanly marked arc and curve systems on $\Omega$.
4.1.1. Witness projections
Let $\Sigma\subset\Omega$ be a compact, essential, non-pants, non-annular subsurface. Let $\mathcal{MS}(\Omega,\Sigma)\subset\mathcal{MS}(\Omega)$ denote the subset of cleanly marked arc and curve systems intersecting $\Sigma$. We construct a projection $\rho_{\Sigma}:\mathcal{MS}(\Omega,\Sigma)\to\mathcal{MS}(\Sigma)$ as follows (see e.g. [Sch, §5.2]). Let $\iota:\Sigma\hookrightarrow\Omega$ be the inclusion map, let $p:\Omega_{\Sigma}\to\Omega$ be the covering space associated to $\pi_{1}(\Sigma)\cong\operatorname{im}\iota_{*}<\pi_{1}(\Omega)$ with Gromov closure $\overline{\Omega}_{\Sigma}$. Let $\tilde{\iota}:\Sigma\hookrightarrow\Omega_{\Sigma}$ be the (unique) lift of $\iota$, and $\bar{\iota}$ its inclusion into $\overline{\Omega}_{\Sigma}$. Fix any homeomorphism $\sigma:\overline{\Omega}_{\Sigma}\to\Sigma$ that is a homotopy inverse for $\bar{\iota}$; note that $\sigma$ is unique up to homotopy, hence isotopy.
Given $\omega\in\mathcal{MS}(\Omega,\Sigma)$, let $\rho_{\Sigma}(\omega)$ be the isotopy class defined by the closures of non-peripheral components of $\sigma p^{-1}(\omega)$. In particular, $\rho_{\Sigma}$ preserves (only) essential curves in $\Sigma$: if $a\in\operatorname{base}\rho_{\Sigma}(\omega)$ is a curve component then $a\in\omega$ and $a$ is essential in $\Sigma$; if $\operatorname{trans}_{\omega}(a)=\pi_{a}b$ then likewise $b\subset\Sigma$ since $\omega$ is cleanly marked and we again assign transversal $\pi_{a}b$.
One verifies that $\rho_{\Sigma}(\omega)$ is cleanly marked and independent of the choice of representative for $\omega$ and $\sigma$. Likewise, $\rho_{\Sigma}$ is independent of the choice of embedding of $\Sigma$: if $\iota^{\prime}:\Sigma\hookrightarrow\Omega$ is isotopic to $\iota$, then the lift $\bar{\iota}\,^{\prime}$ is isotopic to $\bar{\iota}$ and thus a homotopy inverse for $\sigma$.
The natural action of $\operatorname{PMod}(\Sigma)$ on $\mathcal{MS}(\Sigma)$ defines an action of $\operatorname{Mod}(\Sigma,\partial\Sigma)\twoheadrightarrow\operatorname{PMod}(\Sigma)$. Similarly, $\operatorname{Mod}(\Sigma,\partial\Sigma)\curvearrowright\mathcal{MS}(\Omega,\Sigma)$ via the homomorphism $\operatorname{Mod}(\Sigma,\partial\Sigma)\to\operatorname{PMod}(\Omega)$ obtained by extending by identity.
Lemma 4.3.
$\rho_{\Sigma}:\mathcal{MS}(\Omega,\Sigma)\to\mathcal{MS}(\Sigma)$ is $\operatorname{Mod}(\Sigma,\partial\Sigma)$-equivariant.
Proof.
Let $\varphi_{0}\in\operatorname{Mod}(\Sigma,\partial\Sigma)$, fixing a representative. Let $\varphi\in\operatorname{PMod}(\Omega)$ be its extension by identity; since $\varphi$ is (compactly) supported in $\Sigma$, it lifts to a quasi-isometry on $\Omega_{\Sigma}$ that extends to a homeomorphism $\overline{\varphi}$ on $\overline{\Omega}_{\Sigma}$. Since $\overline{\iota}\varphi_{0}=\overline{\varphi}\,\overline{\iota}$ and $\sigma,\overline{\iota}$ are homotopy inverses, $\varphi_{0}\sigma$ and $\sigma\overline{\varphi}$ are homotopic and thus isotopic. For $\omega\in\mathcal{MS}(\Omega,\Sigma)$, $\sigma\overline{p^{-1}(\varphi\omega)}=\sigma\overline{\varphi}\overline{p^{-1}(\omega)}$ is isotopic to $\varphi_{0}\sigma\overline{p^{-1}(\omega)}$, whence the claim follows. ∎
Corollary 4.4.
Let $\phi\in\operatorname{PMod}(\Sigma)$. Then there exists $\psi\in\operatorname{PMod}(\Omega)$ preserving $\mathcal{MS}(\Omega,\Sigma)$ such that for any $\omega\in\mathcal{MS}(\Omega,\Sigma)$, $\phi\rho_{\Sigma}(\omega)=\rho_{\Sigma}(\psi\omega)$. ∎
Given a combinatorial model $\mathcal{M}$ on $\Omega$, let $V(\mathcal{M}),E(\mathcal{M})$ denote its vertex and edge sets, respectively. If $\Sigma$ is a witness for $\mathcal{M}$ then $V(\mathcal{M})\subset\mathcal{MS}(\Omega,\Sigma)$ and $\rho_{\Sigma}$ defines a projection $V(\mathcal{M})\to\mathcal{MS}(\Sigma)$.
Definition 4.5.
A connected combinatorial model $\mathcal{M}$ on $\Omega$ is admissible if
@@ 4.1.1. Witness projections
    $\operatorname{PMod}(\Omega)$ preserves $V(\mathcal{M})$ and extends to an action on $\mathcal{M}$, and
@@
    for any non-annular witness $\Delta\subset\Omega$, there exists $L_{\Delta}$ such that if $(a,b)\in E(\mathcal{M})$, then $i(\rho_{\Delta}(a),\rho_{\Delta}(b))\leq L_{\Delta}$.

### v3

@@ 4.1. Cocompact arc and curve models
An arc and curve model $\mathcal{M}$ on $\Omega$ is a connected simplicial graph with $V(\mathcal{M})\subset\mathcal{K}(\Omega)$ that admits an action of $\operatorname{PMap}_{c}(\Omega)$ induced by the permutation of its vertices. $\mathcal{M}$ is cocompact if this action is cocompact.
Remark 4.2.
If $\Omega$ is finite-type, then (i) $\operatorname{PMap}_{c}(\Omega)=\operatorname{PMap}(\Omega)$and (ii) $\mathcal{M}$is cocompact if and only if $i(u,u)$ and $i(u,v)$ are uniformly bounded for $u\in V(\mathcal{M})$ and $(u,v)\in E(\mathcal{M})$.
Cocompact arc and curve models include most familiar graphs on finite-type surfaces, such as the arc and curve graph $\mathcal{AC}(\Sigma)$, connected subgraphs preserved by $\operatorname{PMap}(\Sigma)$ (e.g. the curve graph and arc graph), and the pants graph. Masur and Minsky’s marking complex $\mathcal{MC}(\Sigma)$ is likewise included, as we will show below.
4.1.1. Cocompact marking graphs
A marking graph $\mathcal{L}$ on a compact surface $\Sigma$ is a connected simplicial graph whose vertices are locally clean markings on $\Sigma$ [Kop23a]. As above, $\mathcal{L}$ is cocompact if $\operatorname{PMap}(\Sigma)$ acts on $\mathcal{L}$ cocompactly by permuting its vertices. We prove that cocompact arc and curve models and cocompact marking graphs are identical, up to $\operatorname{PMap}(\Sigma)$-equivariant quasi-isometry.
For an essential subsurface $W\subset\Sigma$ and $u$ a collection of simple arcs and curves, let $\pi_{W}(u)\subset\mathcal{C}W$ denote union of the (set-valued) projections of the elements in $u$, and let $d_{W}(u,v)\mathrel{\mathop{:}}=\operatorname{diam}(\pi_{W}(u)\cup\pi_{W}(v))$. We show:
Proposition 4.3.
Let $\mathcal{M}$ be a cocompact arc and curve model on a compact surface $\Sigma$. Then there exists a cocompact marking graph $\mathcal{L}_{\mathcal{M}}$ on $\Sigma$ with an identical witness set and an equivariant coarse quasi-isometry $\zeta:\mathcal{M}\to\mathcal{L}_{\mathcal{M}}$ that coarsely preserves projections to witness curve graphs.
Lemma 4.4.
There is a uniform increasing function $\phi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ depending only on $\Sigma$ such that for any $a,b\in\mathcal{AC}(\Sigma)$ and any essential subsurface $W\subset\Sigma$, $\phi(i(a,b))\geq d_{W}(a,b)$.
Proof.
When $a,b$ are both curves, the claim follows from the Choi-Rafi formula for curves [Wat16, Thm. 2.10]. In general, there exists some increasing (affine) function $\phi^{\prime}:\mathbb{R}_{+}\to\mathbb{R}_{+}$ so that $\phi^{\prime}(i(a,b))>i(a^{\prime},b^{\prime})$ for every $a^{\prime}\in\pi_{\Sigma}(a),b^{\prime}\in\pi_{\Sigma}(b)$. We conclude by observing that $\pi_{W}\circ\pi_{\Sigma}$ and $\pi_{W}$ are uniformly boundedly close. ∎
Remark 4.5.
It follows from Remark 4.2 and Lemma 4.4 that for any cocompact arc and curve model $\mathcal{M}$, witness subsurface projections $\pi_{W}:V(\mathcal{M})\to 2^{\mathcal{C}W}$ are uniformly Lipschitz with uniformly bounded vertex images; in general, any subsurface projection has uniformly bounded diameter vertex images.
Definition 4.6.
A generalized marking $\mu=\{(a_{i},t_{i})\}$ on a surface $\Sigma$ is a multicurve $\operatorname{base}\mu=\{a_{i}\}$ along with bounded subsets $t_{i}\subset\mathcal{C}(a_{i})$ for each $a_{i}$. Let $\mu^{\prime}\subset\mu$ be the maximal submarking with only non-empty transversals. If $\operatorname{base}\mu^{\prime}$ is a pants decomposition on $\Sigma\setminus(\mu\setminus\mu^{\prime})$, then $\mu$ is locally complete.
Definition 4.7.
A locally clean marking $\mu^{\prime}=\{(a_{i}^{\prime},t_{i}^{\prime})\}$ is compatible with a generalized marking $\mu=\{(a_{i},t_{i})\}$ if $a_{i}^{\prime}=a_{i}$ (up to reindexing), $t_{i}=\varnothing$ only if $t_{i}^{\prime}=\varnothing$, and $\operatorname{diam}(t_{i}^{\prime},t_{i})$ is minimal among all choices of clean $t_{i}^{\prime}$.
Only a locally complete generalized marking admits compatible locally clean markings. Exactly analogously to [MM00, Lem. 2.8], we have the following for generalized markings:
Lemma 4.8.
For any locally complete generalized marking $\mu=\{(a_{i},t_{i})\}$, there is a constant $n_{0}$ depending only on $\max_{i}(\operatorname{diam}(t_{i}))$ and at least one and at most $n_{0}^{|\mu|}$ compatible locally clean markings. For any compatible marking $\mu^{\prime}=\{(a_{i},t_{i}^{\prime})\}$, $d_{a_{i}}(t_{i},t_{i}^{\prime})<n_{1}$ for a universal constant $n_{1}$.
For each $u\in V(\mathcal{M})$, we first equivariantly construct a corresponding locally clean marking $\mu_{u}$. Choose representatives for $u=\{a_{i}\}$ in pair-wise minimal position; let $\Gamma_{u}=\bigcup_{i}a_{i}$, viewed as a $1$-complex (add basepoints to isolated simple curves as necessary). For a subcomplex $\Gamma^{\prime}\subset\Gamma_{u}$, let $\partial_{\Sigma}(\Gamma^{\prime})$ denote the set of essential, non-peripheral boundary components of a regular neighborhood of $\Gamma^{\prime}\cup\partial\Sigma$, or equivalently the set of non-peripheral boundary components of the essential subsurface filled by $\Gamma^{\prime}$. Fix an enumeration of the edges in $\Gamma_{u}$ and let $\Gamma_{u,j}\subset\Gamma_{u}$ be an exhaustion of $\Gamma_{u}$ by adding the $j$th successive edge; let $F_{u}$ denote the essential subsurface filled by $u$ and $F_{u,j}$ the essential subsurface filled by $\Gamma_{u,j}$. Let $m_{u}$ be the multicurve $\bigcup_{j}\partial_{\Sigma}(\Gamma_{u,j})$. Obtain the generalized marking $\tilde{\mu}_{u}$ by adding to each component $c_{i}\in m_{u}$ the transversal $\pi_{c_{i}}(u)$.
Claim 4.9.
$\tilde{\mu}_{u}$ satisfies the following:
@@ 4.1.1. Cocompact marking graphs
    The submarking $\tilde{\mu}_{u}\setminus\partial F_{u}$ is contained in $F_{u}$ and complete on $F_{u}$.
@@
    $\tilde{\mu}_{u}\cap\partial F_{u}$ has empty transversals, and for all other $(c_{i},t_{i})\in\tilde{\mu}_{u}$, $d_{c_{i}}(t_{i},\pi_{c_{i}}(u))$ is uniformly bounded independently of $u$.
  3. (iii)
    For $c_{i}\in\operatorname{base}\tilde{\mu}_{u}$, $i(c_{i},u)$ is uniformly bounded independently of $u$.
@@
In particular, $\tilde{\mu}_{u}$ is locally complete and any compatible locally clean marking satisfies (i)-(iii).
Proof of claim.
Let $e_{j}=\overline{\Gamma_{u,j}\setminus\Gamma_{u,j-1}}$ be the $j$th edge in $\Gamma_{u}$ and let $\Gamma_{u,0}=\varnothing$. Let $m_{u,k}=\bigcup_{j=1}^{k}\partial_{\Sigma}(\Gamma_{u,j})$ and $m_{u,0}=\varnothing=F_{u,0}$. Inducting on $k\leq|E(\Gamma_{u})|$, we first claim that $m_{u,k}\setminus\partial F_{u,k}$ is a pants decomposition for $F_{u,k}$. In particular, $m_{u}\setminus\partial F_{u}$ is a pants decomposition of $F_{u}$ and (i) follows: since $u$ fills $F_{u}$, it intersects every component $c_{i}$ of $m_{u}\setminus\partial F_{u}$, which then has transversal $\pi_{c_{i}}(u)\neq\varnothing$ in $\tilde{\mu}_{u}$.
By construction $m_{u,k}\setminus\partial F_{u,k}$ is contained in $F_{u,k}$. The claim holds vacuously for $k=0$; suppose it holds for $k-1\geq 0$ and consider the $k$th edge $e_{k}$. Recall that $\partial_{\Sigma}(\Gamma_{u,j})$ is the set of non-peripheral boundary curves in $F_{u,j}$, hence if $F_{u,k}$ and $F_{u,k-1}$ are isotopic then $m_{u,k}=m_{u,k-1}$ and the claim holds by induction. Assume that $F_{u,k}\supsetneq F_{u,k-1}$. If $e_{k}$ is disjoint from $F_{u,k-1}$, then $F_{u,k}$ is obtained from $F_{u,k-1}$ by adding either a disjoint annulus or a disjoint pair of pants; in either case, the claim holds. Else, $F_{u,k-1}$ is obtained by adding a $1$-handle with core $e_{k}$, hence by adjoining a pair of pants along a (non-peripheral) boundary curve $c\in\partial F_{u,k-1}\cap m_{u,k-1}$: $m_{u,k}\setminus\partial F_{u,k}=(m_{u,k-1}\setminus\partial F_{u,k-1})\cup\{c\}$ is a pants decomposition for $F_{u,k}$.
To prove (ii), observe that by construction $\partial F_{u}$ does not intersect $u$, hence $\mu_{u}\cap\partial F_{u}$ has empty transversals; the remaining transversals are uniformly bounded by Remark 4.5. Finally, every component $c\in m_{u}$ is a boundary curve for some $F_{u,k}$, which is filled by $\Gamma_{u,k}\subset\Gamma_{u}$: any essential intersection between $c$ and $u$ implies an essential self-intersection, whence (iii) follows. ∎
Let $\mu_{u}$ be a choice of locally clean marking compatible with $\tilde{\mu}_{u}$. While $\mu_{u}$ and $\tilde{\mu}_{u}$ are non-canonical, we enforce that the choices are $\operatorname{PMap}(\Sigma)$-equivariant: construct $\mu_{u}$ for representatives $u\in V(\mathcal{M})/\operatorname{PMap}(\Sigma)$ and declare $\mu_{\varphi u}\mathrel{\mathop{:}}=\varphi\mu_{u}$ for $\varphi\in\operatorname{PMap}(\Sigma)$.
Proof of Prop. 4.3.
Let $V(\mathcal{L}_{\mathcal{M}})=\{\mu_{u}:u\in V(\mathcal{M})\}$ and obtain $E(\mathcal{L}_{\mathcal{M}})$ by pushing forward the edge relation on $\mathcal{M}$ via $u\mapsto\mu_{u}$; let $\zeta:\mathcal{M}\to\mathcal{L}_{\mathcal{M}}$ be the $\operatorname{PMap}(\Sigma)$-equivariant surjective Lipschitz map induced by $u\mapsto\mu_{u}$. Since $\mathcal{M}$ is cocompact, $\zeta$ implies that $\mathcal{L}_{\mathcal{M}}$ is cocompact.
Let $W\subset\Sigma$ be an essential, non-pants subsurface. Recall that $u$ is filling on $F_{u}$, hence Claim 4.9(i)-(ii) implies $W$ intersects $\mu_{u}$ if and only if it intersects $F_{u}$ if and only if it intersects $u$: the witness set for $\mathcal{L}_{\mathcal{M}}$ is identical to that of $\mathcal{M}$. Claim 4.9(ii) implies that the projection of $u$ and $\mu_{u}$ to annular witnesses parallel to $\mu_{u}$ is uniformly close; similarly, by Claim 4.9(iii) $i(u,\operatorname{base}\mu_{u})$ is uniformly bounded, hence Lemma 4.4 implies that the projection of $u,\mu_{u}$ to all other witnesses is likewise uniformly close.
It remains to show that $\zeta$ is a coarse Lipschitz retraction, hence a quasi-isometry. Since $\mathcal{L}_{\mathcal{M}}$ has finitely many $\operatorname{PMap}(\Sigma)$-orbits of edges, it suffices to show that the fibers of $\zeta$ are uniformly bounded. The following lemma completes the proof. ∎
Lemma 4.10.
The fibers $E_{\mu}=\zeta^{-1}(\mu)$ are uniformly bounded over $\mu\in V(\mathcal{L}_{\mathcal{M}})$.
Proof.
Let $u,v\in E_{\mu}$, hence $F_{u}=F_{v}$ and $\mu\setminus\partial F_{u}$ is complete on $F_{u}$. It suffices to show that $i(u,v)$ is uniformly bounded. Let $u^{\prime}\mathrel{\mathop{:}}=u\setminus\partial F_{u},v^{\prime}\mathrel{\mathop{:}}=v\setminus\partial F_{u}$ denote the subsets of non-boundary elements in $u,v$ respectively. By Claim 4.9(iii) $u^{\prime},v^{\prime}$ intersect each component of $\operatorname{base}\mu$ uniformly many times, independently of $u,v$ and $\mu$, hence $u^{\prime},v^{\prime}$ have at most uniformly many components in each pair of pants $P\subset F_{u}\setminus\mu$. Up to Dehn twists along components in $\operatorname{base}\mu$, these components are among finitely many arcs and curves in $P$; by Claim 4.9(ii) projections to curves in $\operatorname{base}\mu$ are uniformly close to the corresponding transversal in $\mu$ (independently of $u,v,\mu$), hence the order of these twists (and hence the intersection number) for any two components in $P$ is uniformly bounded. It follows that $i(u,v)$ is uniformly bounded independently of $u,v,\mu$. ∎
Proposition 4.11.
The map $\mathcal{M}\mapsto\mathcal{L}_{\mathcal{M}}$ induces a bijection between the $\operatorname{PMap}(\Sigma)$-equivariant quasi-isometry types of cocompact arc and curve models on $\Sigma$ and of cocompact marking graphs on $\Sigma$.
Proof.
Since $\mathcal{M}$ and $\mathcal{L}_{\mathcal{M}}$ are equivariantly quasi-isometric, the induced map is well-defined and injective. We show surjectivity. Given a locally clean marking $\mu=\{(a_{i},t_{i})\}$, where $t_{i}=\varnothing$ or $\pi_{a_{i}}(b_{i})$ for some (unique) clean transverse curve $b_{i}$, let $u_{\mu}=\{a_{i}\}\cup\{b_{i}\}$. Let $\mathcal{L}$ be a cocompact marking graph on $\Sigma$, and let $\mathcal{M}_{\mathcal{L}}$ be the graph obtained as the (connected) push-forward of $\mathcal{L}$ by the map $\psi:\mu\mapsto u_{\mu}$. $\psi$ is $\operatorname{PMap}(\Sigma)$-equivariant, hence $\mathcal{M}_{\mathcal{L}}$ is cocompact. As above, $\psi:\mathcal{L}\to\mathcal{M}_{\mathcal{L}}$ is a surjective Lipschitz retraction, hence a quasi-isometry: any two markings in the fiber $E_{u}=\{\mu:u=u_{\mu}\}$ differ by at most uniformly many flip moves, each of which increases the intersection number by at most $2$. Hence $\mathcal{L}_{\mathcal{M}_{\mathcal{L}}}$ is equivariantly quasi-isometric to $\mathcal{L}$, which suffices. ∎
4.1.2. Geometry in the finite-type case
By [Kop23a, Thm. 2.10], cocompact marking graphs on finite-type surfaces are hierarchically hyperbolic with respect to witness subsurface projection, and this geometry is determined up to equivariant quasi-isometry by the set of connected witnesses [Kop23a, Thm. 2.12]. Let $\mathscr{X}^{\mathcal{M}}$ and $\hat{\mathscr{X}}^{\mathcal{M}}$ denote the sets of witnesses and connected witnesses of an arc and curve model $\mathcal{M}$, respectively. Applying Proposition 4.3, we obtain:
Theorem 4.12.
Let $\mathcal{M}$ be a cocompact arc and curve model on a finite-type hyperbolic, non-pants surface $\Sigma$. Then $(\mathcal{M},\mathscr{X}^{\mathcal{M}})$ is an asymphoric hierarchically hyperbolic space with respect to subsurface projection to witness curve graphs $\pi_{W}:\mathcal{M}\to 2^{\mathcal{C}W},W\in\mathscr{X}^{\mathcal{M}}$.
A (connected) witness set on $\Sigma$ is any collection of essential compact (connected) subsurfaces without pants components that is closed under enlargement and the action of $\operatorname{PMap}(\Sigma)$. From [Kop23a, Thm. 2.12] and Propositions 4.3 and 4.11 we have:
Theorem 4.13.
The map $\mathcal{M}\mapsto\hat{\mathscr{X}}^{\mathcal{M}}$ induces a bijection between coarsely $\operatorname{PMap}(\Sigma)$-equivariant quasi-isometry types of cocompact arc and curve models on $\Sigma$ and connected witness sets on $\Sigma$.
4.2. Witness-cocompactness
In the infinite-type setting, we require a related “local” condition, determined by the existence of projections to cocompact arc and curve models on each witness subsurface. Crucially, these projections admit Lipschitz sections, whence we will obtain asymptotic dimension lower bounds in Section 4.3.
4.2.1. Witness projections
Let $\Sigma\subset\Omega$ be a compact, essential, and connected subsurface. Let $\mathcal{K}(\Omega,\Sigma)\subset\mathcal{K}(\Omega)$ denote the subset of collections of arcs and curves intersecting $\Sigma$. We construct a projection $\rho_{\Sigma}:\mathcal{K}(\Omega,\Sigma)\to\mathcal{K}(\Sigma)$ as follows (see e.g. [Sch, §5.2]). Let $\iota:\Sigma\hookrightarrow\Omega$ be the inclusion map and let $p:\Omega_{\Sigma}\to\Omega$ be the covering space associated to $\pi_{1}(\Sigma)\cong\operatorname{im}\iota_{*}<\pi_{1}(\Omega)$ with Gromov closure $\overline{\Omega}_{\Sigma}$. Let $\tilde{\iota}:\Sigma\hookrightarrow\Omega_{\Sigma}$ be the (unique) lift of $\iota$, and $\bar{\iota}$ its inclusion into $\overline{\Omega}_{\Sigma}$. Fix any homeomorphism $\sigma:\overline{\Omega}_{\Sigma}\to\Sigma$ that is a homotopy inverse for $\bar{\iota}$; note that $\sigma$ is unique up to homotopy, hence isotopy.
Given $\omega\in\mathcal{K}(\Omega,\Sigma)$, let $\rho_{\Sigma}(\omega)$ be the isotopy class defined by the closures of non-peripheral components of $\sigma p^{-1}(a)$ for all $a\in\omega$.
One verifies that $\rho_{\Sigma}(\omega)$ is independent of the choice of representative for $\omega$ and $\sigma$. Likewise, $\rho_{\Sigma}$ is independent of the choice of embedding of $\Sigma$: if $\iota^{\prime}:\Sigma\hookrightarrow\Omega$ is isotopic to $\iota$, then the lift $\bar{\iota}\,^{\prime}$ is isotopic to $\bar{\iota}$ and thus a homotopy inverse for $\sigma$.
The natural action of $\operatorname{PMap}(\Sigma)$ on $\mathcal{K}(\Sigma)$ defines an action of $\operatorname{Map}(\Sigma,\partial\Sigma)\twoheadrightarrow\operatorname{PMap}(\Sigma)$. Similarly, $\operatorname{Map}(\Sigma,\partial\Sigma)\curvearrowright\mathcal{K}(\Omega,\Sigma)$ via the homomorphism $\operatorname{Map}(\Sigma,\partial\Sigma)\to\operatorname{PMap}_{c}(\Omega)$ obtained by extending by identity.
Lemma 4.14.
$\rho_{\Sigma}:\mathcal{K}(\Omega,\Sigma)\to\mathcal{K}(\Sigma)$ is $\operatorname{Map}(\Sigma,\partial\Sigma)$-equivariant.
Proof.
Let $\varphi_{0}\in\operatorname{Map}(\Sigma,\partial\Sigma)$, fixing a representative. Let $\varphi\in\operatorname{PMap}_{c}(\Omega)$ be its extension by identity; since $\varphi$ is (compactly) supported in $\Sigma$, it lifts to a quasi-isometry on $\Omega_{\Sigma}$ that extends to a homeomorphism $\overline{\varphi}$ on $\overline{\Omega}_{\Sigma}$. Since $\overline{\iota}\varphi_{0}=\overline{\varphi}\,\overline{\iota}$ and $\sigma,\overline{\iota}$ are homotopy inverses, $\varphi_{0}\sigma$ and $\sigma\overline{\varphi}$ are homotopic and thus isotopic. For $a\in\omega\in\mathcal{K}(\Omega,\Sigma)$, $\sigma\overline{p^{-1}(\varphi a)}=\sigma\overline{\varphi}\overline{p^{-1}(a)}$ is isotopic to $\varphi_{0}\sigma\overline{p^{-1}(a)}$, whence the claim follows. ∎
Corollary 4.15.
Let $\phi\in\operatorname{PMap}(\Sigma)$. Then there exists $\psi\in\operatorname{PMap}(\Omega)$ preserving $\mathcal{K}(\Omega,\Sigma)$ such that for any $\omega\in\mathcal{K}(\Omega,\Sigma)$, $\phi\rho_{\Sigma}(\omega)=\rho_{\Sigma}(\psi\omega)$. ∎
Given an arc and curve model $\mathcal{M}$ on $\Omega$, let $V(\mathcal{M}),E(\mathcal{M})$ denote its vertex and edge sets, respectively. If $\Sigma$ is a witness for $\mathcal{M}$ then $V(\mathcal{M})\subset\mathcal{K}(\Omega,\Sigma)$ and $\rho_{\Sigma}$ defines a projection $V(\mathcal{M})\to\mathcal{K}(\Sigma)$.
@@ 4.2.1. Witness projections
A connected arc and curve model $\mathcal{M}$ on $\Omega$ is witness-cocompact if
@@
    $\operatorname{PMap}_{c}(\Omega)$ preserves $V(\mathcal{M})$ and extends to an action on $\mathcal{M}$, and
@@
    for any witness $\Delta\subset\Omega$, there exists $L_{\Delta}$ such that if $(a,b)\in E(\mathcal{M})$, then $i(\rho_{\Delta}(a),\rho_{\Delta}(b))\leq L_{\Delta}$.